OpenStudy (anonymous):

A solution of salt and water contains 70 grams of water per 140 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 35 milliliters of the solution. [Round off your answer to two decimal places]

OpenStudy (anonymous):

You might want to put this in the chemistry section bud.

OpenStudy (anonymous):

I will help you out however.

OpenStudy (anonymous):

really? its a math problem though..

OpenStudy (anonymous):

Thanks i would appreciate that.

OpenStudy (anonymous):

We have 70 grams of water/140 mL of solution. 1 mole of water weighs 16 grams. x moles of water would be in 35 mL of solution. Start off by figuring out how grams of water you need for 16 mL of solution. 70 g/140 mL=x g/35 mL x=17.5 g I believe Now that we know we need 8 grams for 16 mL of water, we simply must convert it to moles with some stoichiometry. 17.5 g*(1 mol/16 g)=1.09 moles Notice how I multiplied by what is called the molar mass ratio. The grams canceled out and I am left with moles.

OpenStudy (anonymous):

would you need to round 1.09?

OpenStudy (anonymous):

Isn't that what the question says to do?

OpenStudy (anonymous):

what would u round to